• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name moffit

    Log entry time 16:00:55 on October15,2002

    Entry number 88279

    keyword=Day Shift Summary

    Began shift with 30uA on the 15cm Dummy Cell (1.4x1.4 raster).

    08:22 Left Spectrometer to 39.754deg (2.011GeV)
    09:00 MCC Removes beam for machine developement time.
    09:38 Ed Folts calls and says he's setting the momentum of both arms to zero. A test of the magnets of Hall C are going to trip them anyway.
    11:05 Ed Folts calls and gives the OK to return the spectrometers to their momentum settings.
    11:57 Left arm cycled and momenta reset
    12:57 MD over. Beam returns. 30uA.